# Mutatis Mutandis: Revisiting the Comparator in Discrimination Testing

### v2

@@ 2.3. The Counterfactual Model of Discrimination
Discrimination is often conceived as a causal claim on the (in)direct effect of the protected attribute $A$ on the decision outcome of interest $Y$ (Heckman, 1998). In Example 1.1, e.g., we are interested in whether Clara did not receive tenure because of her gender? Its causal underpinning, which is motivated by non-discrimination law’s own definition of discrimination (Weerts et al., 2023), has long motivated a range of methods for testing discrimination based on counterfactual reasoning. These methods operationalize the scenario in which we are able to manipulate the protected attribute of the individual(s) making the claim, imagine the “what would have been if” (or the counterfactual) outcome, and compare it to the “what is” (or the factual) outcome to isolate the causal effect of $A$ on $Y$. Kohler-Hausmann (2019) refers to this practice as the counterfactual model of discrimination. [^9] Kohler-Hausmann (2019) actually refers to it as the counterfactual causal models of discrimination, which is redundant as counterfactuals are causal (Woodward, 2005). Hence, we drop the “causal.” Using SCM, we can state the discrimination claim as $A\rightarrow Y$ (if direct) or as $A\rightarrow\mathbf{X}\rightarrow Y$ (if indirect). The goal of any discrimination testing tool is validating such claim.
@@
What we want to emphasize about the counterfactual model of discrimination is not that it relies on counterfactual reasoning to answer causal claims, but that, because of this reliance, the comparator $i^{\prime}$ is a counterfactual representation of the complainant $i$. Viewing $i^{\prime}$ as a counterfactual representation of the factual $i$ ultimately positions the comparator as a causal modeling problem. This additional emphasize from our side is important when we consider that $i^{\prime}$ can be either a real or fictitious profile, making it hypothetical in the fullest sense. That the comparator $i^{\prime}$ can be viewed as a thought experiment is not exploited–both conceptually and experimentally–under the CP comparator. We come back to this point in Section 3. We diverge from Kohler-Hausmann (2019) precisely because of this emphasize, viewing the dominant counterfactual model of discrimination as an opportunity suited for fair ML, not as a threat.
@@
Why counterfactuals? We argue, one, because discrimination represents a pattern and, two, because the discrimination claim requires an answer at the individual level. Let us unpack each point. First, why would discrimination by being a pattern amount to a causal relation? Broadly, the presence of a sufficiently robust pattern suggests a causal relationship, as causality is seen as an invariant force acting on the world. [^10] Formally, such view amounts to an interventionist account of causality, which is the basis for both SCM Pearl (2009) and Potential Outcomes Angrist and Pischke (2008), meaning the two most popular causal modeling frameworks. See Woodward (2005) for details. The presence of a causal pattern also suggests, implicitly, that there is little room for chance to explain what we observe. Discrimination claims are often characterized in those terms, meaning an unfair decision that persists over time toward a certain group of individuals. Because of this persistence, we are (somewhat) certain that the unfair treatment comes from the influence of the protected attribute. Intuitively, a discriminatory decision-making process has (un)consciously set up a pattern (i.e., $A\rightarrow Y$ or $A\rightarrow\mathbf{X}\rightarrow Y$) that causes the unfair treatment of protected individuals. If the goal is to detect such pattern(s) and we agree on the premise that discrimination represents a causal pattern, then counterfactual reasoning is the main approach to do so. There is a long tradition in causality, especially within SCM, to formulate what constitutes a cause in terms of counterfactual reasoning (Pearl, 2009; Halpern, 2016; Halpern and Pearl, 2001a, b): if intervening $A$ induces a counterfactual outcome $Y^{CF}$ different from the factual outcome $Y$, then $A$ must be a cause of $Y$ [^11] These definitions are much more formal, such as, in terms of probability distributions, but the phrase captures the main idea behind all of them. Under this setting, to test for the existence of such pattern would imply to examine the impact of $A$ on the decision-making process by changing $A$ itself. If there is such a pattern, then observing a change in outcome due to having changed only $A$ (and controlling for everything else) would, first, confirm the existence of the pattern, and, second, confirm the role of $A$ within it.
@@ 3. Mutatis Mutandis
(The MM Comparator) Let $\tilde{d}$ be a protected-aware distance function that, unlike $d$ in (2), considers both $\mathbf{X}$ and $A$ as inputs, capturing the potential influence of $A$ on $\mathbf{X}$. For the complainant $i$ $\langle\mathbf{x}_{i},a_{i},y_{i}\rangle$, we define a mutatis mutandis (MM) comparator $i^{\prime}$ as an individual profile $\langle\mathbf{x}_{i^{\prime}},a_{i^{\prime}},y_{i^{\prime}}\rangle$ such that:
@@
where ϵ+italic-$\epsilon\in\mathbb{R}^{+}$ is a threshold.
Under $\tilde{d}$, two instances with different non-protected attributes, $\mathbf{x}_{i}\neq\mathbf{x}_{i^{\prime}}$, can be considered similar given information, respectively, about $a_{i}$ and $a_{i^{\prime}}$. The fact that Definition 3.1 allows for this “dissimilarity” between comparator and complainant motivates the MM denomination. The general idea is that under $\tilde{d}$ we should be able, if needed, to account for the influence of $A$ on $\mathbf{X}$ when deriving the $i^{\prime}$ relative to $i$. In other words, we “adjust for what needs to be adjusted” as a consequences on “changing” $A$, allowing us to derive an individual profile less conservative than the one conceived under $d$ in Definition 2.1. We motivate this argument in the next sub-sections.
Given the meaning behind $\tilde{d}$, note that, when $a_{i}=a_{i^{\prime}}$, the MM and CP comparators are the same. As we move within the same group in $A$, we assume that the influence of $A=a$ on $\mathbf{X}$ is shared among all $i$ individuals with $A=a_{i}$. Implicitly, we assume that defining similarity between individuals of the same group in $A$ requires no additional considerations due to shared experience. Additionally, when $\mathbf{X}$ and $A$ are independent, the MM and CP comparators are the same even if $a_{i}\neq a_{a^{\prime}}$ as there are no downstream effects of $A$ to $\mathbf{X}$ to be accounted for.
@@ 3.1. Against Idealized Comparisons
Let us focus on the standard, CP comparator. We argued already that any notion of similarity between individuals implemented through the distance function $d$ in (2) is a normative statement. We also argued that, in turn, the comparator derived under $d$ represents the normative statement. This is true for the standard comparator as for any other formulation of the comparator, such as the MM comparator. Now there are two characteristics specific to the CP comparator worth stressing. First, in Definition 2.1, similarity between individuals comes down to a comparison on the non-protected attributes $\mathbf{X}$ between $i^{\prime}$ and $i$. The distance function $d$ is oblivious to (or unaware of) the protected attribute $A$. This holds whether $A$ and $\mathbf{X}$ are associated or not as $d$ does not take it into consideration. Second, it follows that the counterfactual representation of the complainant $i$ in the form of the comparator $i^{\prime}$ only varies, ideally, in terms of $A$. In other words, nothing would have changed for $i$ in $\mathbf{X}$ had she or he had been from the non-protected group. Because of these two characteristics, we say that the standard CP comparator represents an idealized comparison between $i^{\prime}$ and $i$.
@@
Is the idealized comparison problematic, though? The answer to this questions comes down to having a stance on similarity between individuals, and how it should be operationalized through the $d$ in Definition 2.1. We view it as open to discussion. In this work, we do not have a strong stance against the CP comparator as it can be useful, e.g., in uncovering direct discrimination or when an alternative comparator is not possible. We do, however, view the discussion against it as necessary for discrimination testing. This discussion, e.g., has motivated the MM comparator.
@@
Key in criticizing the CP comparator and the idealized comparison it represents is Kohler-Hausmann (2019). Under the premise that the comparator $i^{\prime}$ in Definition 2.1 represents a counterfactual representation of the complainant $i$, Kohler-Hausmann (2019) criticizes the counterfactual model of discrimination for ignoring how membership to a protected group conditions the material outcomes of individuals and, thus, limits any meaningful counterfactual analysis between protected and non-protected individuals. Kohler-Hausmann (2019), in particular, argues that protected attributes, like race and gender, are social constructs, which is a view widely held by social scientists (Bonilla-Silva, 1997; Sen and Wasow, 2016; Hanna et al., 2020). What it means is that these social categories were once (or still are) used not only to classify individuals but to also delimit the opportunities available to them (Mallon, 2007). According to Kohler-Hausmann (2019), it is inconceivable to picture a counterfactual white version of a black individual as stated in Definition 2.1. Under such definition, the ideal comparator $i^{\prime}$ will have a set of non-protected attributes identical to those of $i$, or $d(\mathbf{x}_{i},\mathbf{x}_{i^{\prime}})=0$. By envisioning a counterfactual world where $i$ can be exactly as he or she is today, in terms of non-protected attributes, while belonging to the non-protect group, this mental exercise requires from us to accept that the protected individual $i$ would have arrived at the same point had he or she been a non-protected individual $i^{\prime}$, which reduces the pervasive influence of $A$.
@@
In short, Kohler-Hausmann (2019) implies that the comparator $i^{\prime}$ derived in Definition 2.1 is wrong and, in turn, the tools for testing discrimination built using such comparator are also wrong. [^12] Kohler-Hausmann (2019) later extends the above argument for gender (Hu and Kohler-Hausmann, 2020), covering the two main protected attributes. We agree partially with her. First, we do not see all discrimination testing that uses the CP comparator as wrong as it might be the only viable options in many cases. Second, we do not picture a better modeling framework for discrimination testing than the counterfactual model of discrimination. It is not ideal, but it can also improve if we are willing to revisit the derivation of the comparator.
To be clear, arguments against discrimination testing are not new. For instance, Heckman (1998) already criticized ideal comparisons in causal modeling. Similarly, Wachter et al. (2020), by classifying fair ML methods as bias preserving or bias transforming depending on how they define similarity, recently pointed out at the limitations of using the conservative CP comparator. These arguments, however, are largely absent from the discussion around discrimination testing tools. There, the focus is always on tool development, which is to be expected. However, as we already argued, the circularity of similarity requires for a deeper considerations from the tool developers.
@@ 3.2. A New Kind of Comparator
Once we allow for a more complex view on the protected attribute $A$ and its influence on the non-protected attributes $\mathbf{X}$, by, e.g., viewing $A$ as a social construct (Kohler-Hausmann, 2019) or acknowledging $A$ as a source of inequality (Michaels and Reed, 2020), the CP comparator appears insufficient. Under this view, there is the need to clarify the relationship between $A$ and $\mathbf{X}$, and whether it is in our interest to account for it when testing for discrimination. The MM comparator, based on $\tilde{d}$ in (4), is meant to clarify and account for the relationship between $A$ and $\mathbf{X}$, embodying what is refereed to as “fairness given the difference” (Álvarez and Ruggieri, 2023). We believe that the MM comparator represents a new kind of comparator, aimed at not only proving discrimination but at shaping the social context behind it.
The MM comparator represents an ideal target level of equality instead of the formal one captured by the CP comparator due its idealized comparison. Conceptually, if $\tilde{d}$ accounts for the influence of $A$ on $\mathbf{X}$, then by comparing $i^{\prime}$ to $i$ when $\mathbf{x}_{i^{\prime}}\neq\mathbf{x}_{i}$ and $a_{i^{\prime}}\neq a_{i}$ we are making a normative statement claiming that these two observably dissimilar individuals should be similar. Under counterfactual reasoning, since the counterfactual question is around $A$, the comparator $i^{\prime}$ derived under Definition 3.1 represents where the complainant would have been if free from the effect of being part of the protected group. The CP comparator does not allow for this possibility as it only envisions a comparator $i^{\prime}$ all equal but for $A$ to the complainant $i$. That is why in Example 1.1, Mike is the CP comparator while Vincent the MM comparator, as Vincent, with his higher number of publications, is meant to represent a Clara that would have carried out her academic career without the social expectations of being the main caregiver of her household.
@@
The MM comparator acts as a target because we, essentially, need to generate it. In Example 1.1, e.g., why should we consider Vincent when other profiles could also meet the MM comparator definition? In other words, we need to have a modeling procedure that motivates the choice of the MM comparator. In practice it means estimating the effect of $A$ on $\mathbf{X}$, and adjusting the latter accordingly. We discuss this last point in the next section. Under this “modeling” interpretation, the proposed MM comparator might seem abstract. In fact, why complicate the derivation of the comparator under the MM definition when we already have the CP definition? A reason is that, as argued by Wachter et al. (2020), non-discrimination law, at least in the EU , aims at substantive equality. In practical terms, it means that when testing for discrimination, we are not interested in whether the treatment was the same for similar individuals, but whether the setting (Wachter et al. (2020) refers to it as the “status quou”) is suitable for similar individuals to be treated similarly over time. This point is similar to the argument on why equal opportunity is better than demographic parity by Hardt et al. (2016): there is a difference between hiring equal number of males and female from ensuring that those numbers remain over time. In fact, this distinction summarizes why Wachter et al. (2020) classifies fair ML definitions into bias preserving (like demographic parity) and bias transforming (like equal opportunity).
@@
A second reason to consider the MM comparator is that, implicitly, we are already implementing methods that aim at ideal targets. We are not proposing a drastic turn in fair ML research. A good example of this point is the fairness-accuracy trade-off (Wick et al., 2019; Sharma et al., 2020; Rodolfa et al., 2021; Maity et al., 2020; Dutta et al., 2020). Recent works on this topic agree that the trade-off exists as long as the data is biased. Hence, demonstrating its existence is trivial given the current state of our society; what is of interest is studying when or under which conditions the trade-off disappears. These are hypothetical conditions that, though, serve as future targets to aim for. Not surprisingly, in a setting in which $\mathbf{X}$ is non-informative of $A$ or, similarly, $A$ does not influence $\mathbf{X}$, there is no need to consider a trade-off between accuracy and fairness. In a fair world, the most accurate model is also the most fair. We believe that the MM comparator is meant to capture such world for the purpose of discrimination testing.
@@ 3.3. Extending the Setup
Consider indirect discrimination cases, meaning $A\rightarrow\mathbf{X}\rightarrow Y$ or $Y=f(\mathbf{X})$. Under the causal modeling problem of deriving a comparator, the key question to be answered is: what happens to $\mathbf{X}$ once we manipulate $A$ to derive the comparator? Given how Definitions 2.1 and 3.1 differ, assuming a case of indirect discrimination, the CP comparator ignores the downstream effect of $A$ on $\mathbf{X}$ while the MM comparator accounts for it. However, while the implementation of the former is clear, how do we actually implement the latter? How do we define $\tilde{d}$ in (4)? Answering this question helps extend the discrimination testing setup in Section 2.1. To start, let $g$ represent a generative model such that:
(5) $\text{for}\;\;\tilde{X}=g(\mathbf{X},A)\;\;\;\text{we have}\;\;\;\tilde{X}\perp\!\!\!\!\perp A$
We can view, broadly, $\tilde{X}$ as an adjusted representation of $X$ that is removed from the influence of $A$. Conceptually, (5) describes the problem of fair representation learning (Zemel et al., 2013) in which we want to learn or have access to $g$ in order to achieve $\tilde{X}$. Such a model, e.g., can be a Variational Auto-Encoder or a Markov Chain (Kingma and Welling, 2014; Salimans et al., 2015; Bengio et al., 2013).
For our purposes, we view $\tilde{X}$ as a counterfactual (fair) representation of $\mathbf{X}$ that answers to the question of what would have happened had the individual belonged to the non-protected group instead of the protected group? Hence, $g$ represents the counterfactual generation procedure described in Section 1.1. Given $g$, we should be able to generate the set of non-protected attributes of the complainant $i$’s comparator $i^{\prime}$ and derive the MM comparator $i^{\prime}$ such that:
@@
where, possibly, $\tilde{\mathbf{x}}_{i}\neq\mathbf{x}_{i}$. Given the causal relation $A\rightarrow\mathbf{X}$, the comparator $i^{\prime}$ is similar to the counterfactual profile of $i$ but dissimilar to its factual profile. Here, $d$ represents again a distance function unaware of the protected attribute $A$ in (2). Through (6), we summarize the link between the Definitions 2.1 and 3.1 complainant $i$ and its comparator $i^{\prime}$ as:
@@
where the left-hand-side of the equation represents the general definition of the MM comparator while the right-hand-side an implementation of it under the generative model $g$. The relationship (7) allows to derive the MM comparator via $\tilde{d}$ or via $g$ plus $d$. Our preference, based on our counterfactual reasoning approach, is with using $g$ plus $d$, which allows for: (i) making it clear the difference between the MM and the CP approaches; (ii) reusing the existing notions of distances and algorithms for inferring them (Kulis, 2013; Ilvento, 2020). In that case, the setup in Section 2.1 would require an additional element on top of deriving the comparator and comparing the outcomes: learning $g$, which is of course not limited to counterfactual generation. This extended setup for discrimination testing represents an opportunity for ML generative models and a promising path for future work.

### v3

@@ 2.3. The Counterfactual Model of Discrimination
Discrimination is often conceived as a causal claim on the (in)direct effect of the protected attribute $A$ on the decision outcome of interest $Y$ (Heckman, 1998). Its causal underpinning, which is motivated by non-discrimination law’s own definition of discrimination (Weerts et al., 2023), has long motivated a range of methods for testing discrimination based on counterfactual reasoning. These methods operationalize the scenario in which we are able to manipulate the protected attribute of the individual(s) making the claim, imagine the “what would have been if” (or the counterfactual) outcome, and compare it to the “what is” (or the factual) outcome to isolate the causal effect of $A$ on $Y$. Kohler-Hausmann (2019) refers to this practice as the counterfactual model of discrimination. [^10] Kohler-Hausmann (2019) actually refers to it as the counterfactual causal models of discrimination, which is redundant as counterfactuals are causal (Woodward, 2005). Hence, we drop the “causal.” Using SCM, we can state the discrimination claim as $A\rightarrow Y$ (if direct) or as $A\rightarrow\mathbf{X}\rightarrow Y$ (if indirect). The goal of any discrimination testing tool is validating such claim.
@@
What we want to emphasize about the counterfactual model of discrimination is not that it relies on counterfactual reasoning to answer causal claims, but that, because of this reliance, the comparator $i^{\prime}$ is a counterfactual representation of the complainant $i$. Viewing $i^{\prime}$ as a counterfactual representation of the factual $i$ ultimately positions the comparator as a causal modeling problem. This additional emphasize from our side is important when we consider that $i^{\prime}$ can be either a real or fictitious profile, making it hypothetical in the fullest sense. Further, that the comparator $i^{\prime}$ can be viewed as a thought experiment is not exploited—both conceptually and experimentally—under the CP comparator. We come back to this point in Section 3. This work diverges from Kohler-Hausmann (2019) precisely because of this emphasize, viewing the dominant counterfactual model of discrimination as an opportunity suited for fair ML.
@@
Why counterfactuals? We argue, one, because discrimination represents a pattern and, two, because the discrimination claim requires an answer at the individual level. Let us unpack each point. First, why would discrimination by being a pattern amount to a causal relation? Broadly, the presence of a sufficiently robust pattern suggests a causal relationship, as causality is seen as an invariant force acting on the world. [^11] Formally, such view amounts to an interventionist account of causality, which is the basis for both SCM (Pearl (2009)) and Potential Outcomes (Angrist and Pischke (2008)), meaning the two most popular causal modeling frameworks. See Woodward (2005) for details. The presence of a causal pattern also suggests, implicitly, that there is little room for chance to explain what we observe. Discrimination claims are often characterized in those terms, meaning an unfair decision that persists over time toward a certain group of individuals. Because of this persistence, we are (somewhat) certain that the unfair treatment comes from the influence of the protected attribute. Intuitively, a discriminatory decision-making process has (un)consciously set up a pattern (i.e., $A\rightarrow Y$ or $A\rightarrow\mathbf{X}\rightarrow Y$) that causes the unfair treatment of protected individuals. If the goal is to detect such pattern(s) and we agree on the premise that discrimination represents a causal pattern, then counterfactual reasoning is the main approach to do so. There is a long tradition in causality, especially within SCM, to formulate what constitutes a cause in terms of counterfactual reasoning (Pearl, 2009; Halpern, 2016; Halpern and Pearl, 2001a, b): if intervening $A$ induces a counterfactual outcome $Y^{CF}$ different from the factual outcome $Y$, then $A$ must be a cause of YY. [^12] These definitions are much more formal, such as, in terms of probability distributions, but the phrase captures the main idea behind all of them. Under this setting, to test for the existence of such pattern would imply to examine the impact of $A$ on the decision-making process by changing $A$ itself. If there is such a pattern, then observing a change in outcome due to having changed only $A$ (and controlling for everything else) would, first, confirm the existence of the pattern, and, second, confirm the role of $A$ within it.
@@ 3. Mutatis Mutandis
[…]
@@
where $\epsilon\in\mathbb{R}^{+}$ is a threshold. Recall that $i$ must be real profile; $i^{\prime}$ can instead be real or fictitious.
[…]
@@ 3.1. Against Idealized Comparisons
Let us focus on the standard CP comparator. We argued already that any notion of similarity between individuals implemented through the distance function $d(\cdot,\cdot)$ in Equation (2) is a normative statement. We also argued that, in turn, the comparator derived under $d(\cdot,\cdot)$ represents the normative statement. This is true for the standard comparator as for any other formulation of the comparator, such as the MM comparator. Now there are two characteristics specific to the CP comparator worth stressing. First, in Definition 2.1, similarity between individuals comes down to a comparison on the non-protected attributes $\mathbf{X}$ between $i^{\prime}$ and $i$. The distance function $d(\cdot,\cdot)$ is oblivious to (or unaware of) the protected attribute $A$. This holds whether $A$ and $\mathbf{X}$ are associated or not as $d(\cdot,\cdot)$ does not take it into consideration. Second, it follows that the counterfactual representation of the complainant $i$ in the form of the comparator $i^{\prime}$ only varies, ideally, in terms of $A$. In other words, nothing would have changed for $i$ in $\mathbf{X}$ had she or he had been from the non-protected group. Because of these two characteristics, we say that the standard CP comparator represents an idealized comparison between $i^{\prime}$ and $i$.
@@
Is the idealized comparison problematic, though? The answer to this questions comes down to having a stance on similarity between individuals, and how it should be operationalized through the $d(\cdot,\cdot)$ in Definition 2.1. We view it as open to discussion. In this work, we do not have a strong stance against the CP comparator as it can be useful, for example, in uncovering direct discrimination or when an alternative comparator is not possible. We do, however, view the discussion against it as necessary for discrimination testing. This discussion motivated the MM comparator.
@@
Pivotal in criticizing the CP comparator and the idealized comparison it represents is Kohler-Hausmann (2019). Under the premise that the comparator $i^{\prime}$ in Definition 2.1 represents a counterfactual representation of the complainant $i$, Kohler-Hausmann (2019) criticizes the counterfactual model of discrimination for ignoring how membership to a protected group conditions the material outcomes of individuals and, thus, limits any meaningful counterfactual analysis between protected and non-protected individuals. Kohler-Hausmann (2019), in particular, argues that protected attributes, like race and gender, are social constructs, which is a view widely held by social scientists (Bonilla-Silva, 1997; Sen and Wasow, 2016; Hanna et al., 2020). What it means is that these social categories were once (or still are) used not only to classify individuals but to also delimit the opportunities available to them (Mallon, 2007). According to Kohler-Hausmann (2019), it is inconceivable to picture a counterfactual white version of a black individual as stated in Definition 2.1. Under such definition, the ideal comparator $i^{\prime}$ will have a set of non-protected attributes identical to those of $i$, or $d(\mathbf{x}_{i},\mathbf{x}_{i^{\prime}})=0$. By envisioning a counterfactual world where $i$ can be exactly as he or she is today, in terms of non-protected attributes, while belonging to the non-protect group, this mental exercise requires from us to accept that the protected individual $i$ would have arrived at the same point had he or she been a non-protected individual $i^{\prime}$, which reduces the pervasive influence of $A$.
@@
Kohler-Hausmann (2019) implies that the comparator $i^{\prime}$ derived in Definition 2.1 is wrong and, in turn, the tools for testing discrimination built using such comparator are also wrong. [^13] Kohler-Hausmann (2019) later extends the above argument for gender (Hu and Kohler-Hausmann, 2020), covering the two main protected attributes. We agree partially with her. First, we do not see all discrimination testing that uses the CP comparator as wrong as it might be the only viable options in many cases. Second, we do not picture a better modeling framework for discrimination testing than the counterfactual model of discrimination. It is not ideal, but it can also improve if we are willing to revisit the derivation of the comparator. In Appendix A we discuss additional critical work toward discrimination testing from a non-legal angle.
Arguments against discrimination testing are not new. Similarly, Wachter et al. (2020), by classifying fair ML methods as bias preserving or bias transforming depending on how they define similarity, recently pointed out at the limitations of using the conservative CP comparator. These arguments, however, are largely absent from the discussion around discrimination testing tools. There, the focus is always on tool development, which is to be expected. However, as we already argued, the circularity of similarity requires for a deeper considerations from the tool developers.
@@ 3.2. A New Kind of Comparator
[…]
@@
The MM comparator acts as a target because we, essentially, need to generate it. In Example 1.1, for example, why should we consider Vincent when other profiles could also meet the MM comparator definition? In other words, we need to have a modeling procedure that motivates the choice of the MM comparator. In practice it means estimating the effect of $A$ on $\mathbf{X}$, and adjusting the latter accordingly. We discuss this last point in the next section. Under this “modeling” interpretation, the proposed MM comparator might seem abstract. In fact, why complicate the derivation of the comparator under the MM definition when we already have the CP definition? A reason is that, as argued by Wachter et al. (2020), non-discrimination law, at least in the EU , aims at substantive equality. In practical terms, it means that when testing for discrimination, we are not interested in whether the treatment was the same for similar individuals, but whether the setting (Wachter et al. (2020) refers to it as the “status quou”) is suitable for similar individuals to be treated similarly over time. This point is similar to the argument on why equal opportunity is better than demographic parity by Hardt et al. (2016): there is a difference between hiring equal number of males and female from ensuring that those numbers remain over time. In fact, this distinction summarizes why Wachter et al. (2020) classifies fair ML definitions into bias preserving (like demographic parity) and bias transforming (like equal opportunity).
@@
A second reason to consider the MM comparator is that, implicitly, we are already implementing methods that aim at ideal targets. We are not proposing a drastic turn in fair ML research. A good example of this point is the fairness-accuracy trade-off discussion (Wick et al., 2019; Sharma et al., 2020; Rodolfa et al., 2021; Maity et al., 2020; Dutta et al., 2020). Recent works on this topic agree that the trade-off exists as long as the data is biased. Hence, demonstrating its existence is trivial given the current state of our society; what is of interest is studying when or under which conditions the trade-off disappears. These are hypothetical conditions that, though, serve as future targets to aim for. Not surprisingly, in a setting in which $\mathbf{X}$ is non-informative of $A$ or, similarly, $A$ does not influence $\mathbf{X}$, there is no need to consider a trade-off between accuracy and fairness. In a fair world, the most accurate model is also the most fair. We believe that the MM comparator is meant to capture such world for the purpose of discrimination testing.
@@ 3.3. Extending the Setup
Let us consider indirect discrimination cases, meaning $A\rightarrow\mathbf{X}\rightarrow Y$. Under the causal modeling problem of deriving a comparator, the key question to be answered is: what happens to $\mathbf{X}$ once we manipulate $A$ to derive the comparator? Given how Definitions 2.1 and 3.1 differ, in the case of indirect discrimination the CP comparator ignores the downstream effect of $A$ on $\mathbf{X}$ while the MM comparator accounts for it. However, while the implementation of the CP comparator is clear, how do we implement the MM comparator? Answering this question helps extend the discrimination testing setup in Section 2.1. To start, let $g(\cdot)$ represent a generative model such that:
(5) $\text{for}\;\;\tilde{\mathbf{X}}=g(\mathbf{X},A)\;\;\;\text{we have}\;\;\;\tilde{\mathbf{X}}\perp\!\!\!\!\perp A$
We view $\tilde{X}$ as an adjusted representation of $\mathbf{X}$ that is removed from the influence of the protected attribute $A$. Conceptually, Equation (5) describes the problem of fair representation learning (Zemel et al., 2013) in which we want to learn or have access to $g(\cdot)$ to achieve $\tilde{\mathbf{X}}$. Such a model, for example, can be a Variational Auto-Encoder or a Markov Chain (Kingma and Welling, 2014; Salimans et al., 2015; Bengio et al., 2013).
For our purposes, we view $\tilde{\mathbf{X}}$ as a counterfactual (fair) representation of $\mathbf{X}$ that answers to the question of what would have happened had the individual belonged to the non-protected group instead of the protected group? Hence, $g(\cdot)$ represents the counterfactual generation procedure of abduction, action, and prediction described in Section 1.1. Given $g(\cdot)$, we should be able to generate the set of non-protected attributes of the complainant $i$’s comparator $i^{\prime}$ and derive the MM comparator $i^{\prime}$ such that:
@@
where, possibly, $\tilde{\mathbf{x}}_{i}\neq\mathbf{x}_{i}$. Given the causal relation $A\rightarrow\mathbf{X}$, the comparator $i^{\prime}$ is similar to the counterfactual profile of $i$ but dissimilar to its factual profile. Here, $d(\cdot,\cdot)$ represents again a distance function unaware of the protected attribute $A$ in Equation (2). Through Equation (6), we summarize the link between the Definitions 2.1 and 3.1 complainant $i$ and its comparator $i^{\prime}$ as:
@@
where the left-hand-side of the equation represents the general definition of the MM comparator while the right-hand-side an implementation of it under the generative model $g(\cdot)$.
[…]